During 2016 and 2017, Faulkner Manufacturing used the sum-of-the-years’-digits (SYD) method of depreciation for its depreciable assets, for both financial reporting and tax purposes. At the beginning of 2018, Faulkner decided to change to the straight-line method for both financial reporting and tax purposes. A tax rate of 40% is in effect for all years.

For an asset that cost $26,200 with an estimated residual value of $1,200 and an estimated useful life of 10 years, the depreciation under different methods is as follows:

Year Straight Line SYD Difference
2016 $ 2,500 $ 4,545 $ 2,045
2017 2,500 4,091 1,591
$ 5,000 $ 8,636 $ 3,636


Required:
1. Prepare the journal entry that Faulkner will record in 2018 related to the change.
2. Suppose instead that Faulkner previously used straight-line depreciation and changed to sum-of-the-years’- digits in 2018. Prepare the journal entry that Faulkner will record in 2018 related to the change.
Answer #1

Answer -

1. Answer -

Event Accounts Titles Debit Credit
1. Depreciation expense $2045.50
   Accumulated depreciation $2045.50

Calculation:

The change in depreciation method from SYD to straight-line is considered a change in estimate and handled prospectively.

In 2018, the remaining book value = $26200 - $8636 = $17564

Remaining life = 8 years (10 – 2)

Residual value = $1200

Straight-line depreciation = ($17564 - $1200) / 8 = $2045.5

A footnote disclosure note should justify that the change is preferable and describe the effect of the change on any financial statement line items and earning per share amounts affected for all periods reported.

Note: Any journal entry Faulkner will record in 2018 related to the change and any required footnote disclosure.

2. Answer -

Event Accounts Titles Debit Credit
2. Depreciation expense $4444
   Accumulated depreciation $4444

Calculation:

Suppose instead that Faulkner previously used straight-line depreciation and changed to SYD method in 2018.

In 2018, the remaining book value = $26200 - $5000 = $21200

Remaining life = 8 years (10 – 2)

Residual value = $1200

The SYD method:

Depreciation expense = Value to depreciate * Depreciation coefficient

Depreciation coefficient = Decreasing year / Digits sum

The digits sum is the sum of all the useful life years, in this case is:

Digits sum = 1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 = 36

Depreciation coefficient = 8 / 36 = 0.2222

SYD depreciation = ($21200 - $1200) / 0.222 = $4444

A footnote disclosure note should justify that the change is preferable and describe the effect of the change on any financial statement line items and earning per share amounts affected for all periods reported.

Note: Any journal entry Faulkner will record in 2018 related to the change and any required footnote disclosure.